OpenStudy (genny7):

Help! Please! -8-4-(-2)-(+2)(-3)

OpenStudy (mrnood):

first you must put any 'minus' signs inside the brackets they refer to. Rewrite the sum with the brackets evaluated...

OpenStudy (mrnood):

(btw - have you missed out a + or - between the last 2 brackets?)

OpenStudy (genny7):

This is order of operations

OpenStudy (mrnood):

and first is B for brackets so do what I suggested first

OpenStudy (genny7):

@MrNood Im stuck

OpenStudy (anonymous):

It can either be PEDMAS or BEDMAS P = Parentheses B = Bracket

OpenStudy (anonymous):

Parentheses, Exponents, Multiplication, Division, Addition, Subtraction

OpenStudy (genny7):

Okay but i keep getting -16 & in my book it has a different answer

OpenStudy (anonymous):

-8-4-(-2)-(+2)(-3) so using the rules we see that ((+2)(-3) would come first

OpenStudy (genny7):

Yes and that would be -6 right

OpenStudy (anonymous):

so then you have -8-4-(-2)-(-6)

OpenStudy (anonymous):

right now do the multiplication of -1*(-2) and -1*(-6)

OpenStudy (anonymous):

so that should give you -8-4+2+6

OpenStudy (anonymous):

from there you can just do -8-4 = -12 and 2+6 = 8 and add the results together

OpenStudy (genny7):

Okay i think i got it from there

OpenStudy (mathstudent55):

This is an "order of operations" problem. In this case, the parentheses are here for only two reasons: 1) to separate a negative sign from an addition sign 2) to show a product Using PEMDAS, you can skip PE, and go right to MDAS which means the following: MD - multiplications and divisions in the order they appear from left to right AS - additions and subtractions in the order they appear from left to right \(-8-4-(-2)-(+2)(-3)\)

OpenStudy (mathstudent55):

\(-8-4-(-2)-\color{red}{(+2)(-3)}\) The operation shown in read above is a multiplication, so it has to be done first. \(= -8-4-(-2)-\color{red}{(-6)}\) Now all you have left is subtractions, so do them in order from left to right, one at a time. \(= -8-4-(-2)-(-6)\) \(= -12-(-2)-(-6)\) \(= -10-(-6)\) \(= -4\)